# DriveMoE: Mixture-of-Experts for Vision-Language-Action Model in End-to-End Autonomous Driving

### v1

@@ 5 Conclusion
In this paper, we introduced DriveMoE improving from our Dirve-$\pi_{0}$, a novel end-to-end autonomous driving framework that integrates Mixture-of-Experts (MoE) architectures into both vision and action components. DriveMoE effectively addresses challenges inherent in existing VLA models by dynamically selecting relevant camera views through a Scene-Specialized Vision MoE, and by employing a Skill-Specialized Action MoE that activates expert modules tailored to specific driving behaviors. Extensive evaluations on the Bench2Drive benchmark demonstrate that DriveMoE achieves state-of-the-art performance, significantly enhancing computational efficiency and robustness to rare, safety-critical driving scenarios. The introduction of MoE into end-to-end driving opens promising avenues for future research, and we will publicly release our code and models to facilitate continued exploration and advancement in this domain.
Appendix A Annotation for Router
@@
Vision Router: We developed a set of heuristic rules based on annotation information from the Bench2Drive dataset to identify special driving scenarios, enabling effective camera-view-level supervision. The Camera Annotation Rules are,
@@
Action Router: As shown in Table 7, Bench2Drive jia2024bench divides 44 scenarios into 5 skills.
@@
Appendix B Limitations and Social Impact
Limitations: DriveMoE is the first end-to-end autonomous driving method to integrate Mixture-of-Experts (MoE) architectures within both vision and action components. Although DriveMoE demonstrates superior performance in empirical evaluations, effectively achieving load balancing among experts remains a significant challenge as the number of experts grows. Future research directions may include exploring adaptive expert assignment and dynamic routing strategies, which could enhance computational efficiency and scalability, ultimately improving the generalization and industrial applicability of end-to-end autonomous driving solutions.
Social Impact: DriveMoE introduces an efficient and effective Mixture-of-Experts-based VLA framework for end-to-end autonomous driving, addressing inefficiencies in multi-view processing and diverse driving behaviors. DriveMoE has significant potential for practical application in industry due to its simplicity and efficiency.
Appendix C Visualization
See the supplementary material DriveMoE.mov file for details.

### v2

@@ 4 Conclusion
We propose DriveMoE, a novel end-to-end autonomous driving framework built upon Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$, which integrates Mixture-of-Experts(MoE) into both vision and action components. DriveMoE effectively addresses challenges inherent in existing VLA models by dynamically selecting relevant camera views through a Scene-Specialized Vision MoE, and by employing a Skill-Specialized Action MoE that activates expert modules tailored to specific driving behaviors. Extensive evaluations on the Bench2Drive benchmark show that DriveMoE achieves state-of-the-art performance, significantly enhancing computational efficiency and robustness to rare, safety-critical driving scenarios.
Supplementary Material
5 Related Work
5.1 VLM/VLA in End-to-end Autonomous Driving
The advancement of Large Language Models (LLMs) has significantly accelerated the development of Vision-Language Models (VLMs) for autonomous driving. Leveraging powerful generalization, open-set reasoning, and scalability, these models have become influential paradigms for end-to-end driving tasks.
DriveGPT-4 [56] pioneers the integration of multimodal LLMs into end-to-end driving by introducing a vision–language–action framework that jointly performs control prediction and natural-language explanation. DriveLM [47] further explores the reasoning aspect of driving intelligence through graph visual question answering (GVQA). LMDrive [45] formulates perception and planning tasks as sequences of discrete tokens, enabling better interpretability and facilitating cross-domain knowledge transfer. LeapAD [40] adopts a data-driven and knowledge-driven Vision-Language Model (VLM) approach. SimLingo [43] proposes a fully closed-loop VLA architecture aligning language reasoning with low-level control. To enhance learning efficiency, it introduces a data-bucket sampling scheme that focuses on diverse and high-risk scenarios, avoiding overfitting to trivial straight-driving cases.
However, Existing approaches predominantly rely on discrete token-based for driving policy learning, without exploring continuous tokenization or diffusion-based policies that could bridge vision-language understanding and continuous control. Moreover, despite the increasing interest in multimodal skill modeling for complex driving scenes, no prior work has investigated Mixture-of-Experts (MoE) architectures to enhance specialization and generalization across diverse driving domains.
To address this limitation, the embodied AI community has proposed vision-language-action (VLA) models that represent actions as continuous variables instead of discrete tokens. Methods such as OpenVLA [33], Diffusion Policy [6] and $\pi_{0}$ [3] demonstrate strong performance by modeling continuous action distributions through sequence prediction and global optimization. Nevertheless, these approaches often rely on task-specific policies or instruction-conditioned models, which struggle to generalize across the long-tail distribution of behaviors seen in complex driving environments.
5.2 Mixture-of-Experts in Large Language Models
Recent progress in Mixture-of-Experts (MoE) architectures has demonstrated remarkable efficiency and scalability in large language models (LLMs). Sparse Mixture-of-Experts (MoE) architectures have become a mainstream approach for scaling LLMs. By replacing the standard feedforward layers in Transformers with expert modules, models like DeepSeekMoE [8] and Mixtral-8x7B [8] improve task specialization and representation capacity while maintaining inference efficiency through conditional computation. In robotics, MoE architectures have also been used to address task heterogeneity and long-tailed data distributions. For example, MENTOR [17] replaces the MLP backbone with MoE layers to enable gradient routing among modular experts, helping mitigate gradient interference in multi-task learning. Despite promising results in language modeling and robot policy learning, the use of MoE in end-to-end autonomous driving remains underexplored.
Motivated by these gaps, our metheds is the first end-to-end autonomous driving framework that integrates MoE at both the vision and action levels.
6 Conditional Flow Matching Loss
Following prior work [3, 37, 38], our method predicts future action trajectories in a denoising manner using a conditional flow matching loss,
$\small L^{\tau}(\theta)=\mathbb{E}_{p(\mathbf{A}_{t}\mid\mathbf{o}_{t}),\,q(\mathbf{A}_{t}^{\tau}\mid\mathbf{A}_{t})}\left\|\mathbf{v}_{\theta}(\mathbf{A}_{t}^{\tau},\mathbf{o}_{t})-\mathbf{u}(\mathbf{A}_{t}^{\tau}\mid\mathbf{A}_{t})\right\|^{2}$ (7)
where subscripts denote timesteps and superscripts denote flow matching timesteps, with $\tau\in[0,1]$. We sample noisy actions $\mathbf{A}_{t}^{\tau}=\tau\mathbf{A}t+(1-\tau)\epsilon$ and train the network to output a denoising flow $\mathbf{v}_{\theta}(\mathbf{A}_{t}^{\tau},o_{t})$ that matches the ground-truth direction $\mathbf{u}(\mathbf{A}_{t}^{\tau}\mid\mathbf{A}_{t})=\epsilon-\mathbf{A}_{t}$. This formulation enables the model to learn the underlying trajectory distribution by aligning with a continuous stochastic process, rather than relying on pointwise supervision. It is particularly suitable for multimodal or uncertain planning scenarios in autonomous driving, where modeling smooth action trajectories is critical.
7 Implementation Details
Vision Routing Annotations: We introduce additional camera-view importance annotations into the Bench2Drive [27] dataset. This annotation approach is both inexpensive and straightforward, yet it significantly improves model performance through efficient and effective utilization of multi-camera inputs. The details about camera annotation rules refer to Appendix 9.
Action Routing Annotations: We maintain skill definitions consistent with Bench2Drive [27] setup. There are five driving skills: Merging, Overtaking, Emergency Brake, Give Way, and Traffic Sign.
Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$: We utilize 2 sequential front-view images as input to our model to effectively estimate the velocities of surrounding traffic agents. Additionally, the input state incorporates both current and historical information, including position, velocity, acceleration, and heading angle, enabling the model to predict 10 future waypoints accurately.
Training Stage 1. We train the model for 12 epochs. The Vision-Language Model (VLM) component is initialized from the pretrained weights of Paligemma-3b-pt-224 [2]. The VLM and Action MoE experts are optimized separately using two optimizers, both configured as follows: learning rate = $5\times 10^{-5}$, and warmup steps enabled. Gradient clipping is applied with a maximum gradient norm of 1.0. Gradient accumulation is used to simulate a batch size of 128. To balance different loss components effectively, we set the vision router loss weight $\lambda_{0}$ to 10.0, action router loss weight $\lambda_{2}$ to 10.0, flow matching loss weight $\lambda_{1}$ to 1.0.
Training Stage 2. We continue training for an additional 6 epochs, initializing from the checkpoint obtained at the end of Stage 1. In this stage, input camera views and action experts are dynamically selected based on outputs from the routers.the vision router loss weight $\lambda_{0}$ to 5.0, action router loss weight $\lambda_{2}$ to 5.0, flow matching loss weight $\lambda_{1}$ to 1.0, emphasizing trajectory learning. Other hyperparameters remain consistent with Stage 1.
8 Discussion on Unsupervised View Selection and Knowledge Distillation
About unsupervised view selection, unsupervised camera view learning could further reduce annotation costs and improve generalization. Our supervised annotations are inexpensive (simple heuristics from trajectory/map) and provide stable training for initial Vision MoE exploration. Nevertheless, the methods DySS [59] and LightVLA [32] offer valuable insights that could inspire image-level pruning. Potential extensions include image-level token pruning after view selection.
About Knowledge Distillation, Combining DriveMoE with knowledge distillation [15] is a promising direction—our architecture could serve as a teacher model to distill skill-specific compact students for deployment.
9 Annotation for Router
@@
Vision Router: We developed a set of heuristic rules based on annotation information from the Bench2Drive dataset to identify special driving scenarios, enabling effective camera-view-level supervision. We select camera views contextually, defaulting to the rear view if no critical view is identified. The Camera Annotation Rules are,
@@
Action Router: As shown in Table 8, Bench2Drive [27] divides 44 scenarios into 5 skills.
@@
10 Experiment on nuScenes
Table 9: Open-loop planning performance in nuScenes
[…]
1s 2s 3s Avg. 1s 2s 3s Avg.
[…]
Following the same skill definitions as Bench2Drive, DriveMoE achieves competitive L2 error while significantly reducing collision rate (Table 9), demonstrating strong generalization to real-world scenarios. Same to [12], open-loop evaluation cannot reflect driving performance.
11 Hyperparameters & Efficienc
As shown in Table 10, the increased cost primarily stems from processing one additional camera view and the MoE modules with Top-3 expert activation. We will clarify this in our revision.
Table 10: Comparison of model scale and inference cost.
[…]
Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ 2 Fixed 63.26 31.82 2606 3400 240 80 GPU-hours
Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ 6 Fixed 62.27 31.36 2606 7576 700 320 GPU-hours
DriveMoE 2 Dynamic 74.22 48.64 3008 3896 260 120/80 GPU-hours